Dinesh S. Thakur

Education: School: Balmohan Vidyamandir, Bombay. College: Ruia College, Bombay. B. Sc. (Math). 1981 University of Bombay. M.A., Ph. D. 1987 Harvard University.

Employment: 1987-1989 Institute for Advanced Study, Princeton. 1989-1991 Tata Institute of Fundamental Research, Bombay. 1991-1992 University of Minnesota. 1992-1993 University of Michigan. 1993- University of Arizona. Secondary: 98-99 TIFR, Mumbai and 2000-2001 IAS, Princeton.

Personal data Born in Mumbai (Bombay), India 1961. Citizen of USA with the dual citizenship of India. Wife: Jyoti. Son: Ashutosh.